# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

### v2

@@ 5.2 Cross-graph and Cross-task Transferability (RQ2)
Setting Cora Citeseer PubMed
Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
zero-shot 18.54 12.16 27.64 17.10 39.12 39.84
@@
5.3 Exploration Studies
Vocabulary Analysis (RQ3). We evaluate the quality of our graph vocabulary using two metrics in Table 8. (a) Coverage. We compute the proportion of nodes that can be matched to IDs. Since the vocabulary is language-based, any node’s semantics can, in principle, be expressed through a combination of language tokens. Thus, every node is successfully assigned a valid ID, confirming complete coverage. (b) Uniqueness. We measure the entropy of token distributions across nodes in both individual and merged graphs. Higher entropy indicates more diverse and fine-grained token usage, indicating stronger uniqueness and representational capacity. Notably, the merged vocabulary (“ALL”) achieves the highest entropy (12.2761), showing that integrating multiple graphs improves vocabulary expressiveness and discriminability, which reinforces the necessity of building GFMs.
@@ 5.3 Exploration Studies
Ablation Studies (RQ4). We construct three ablated variants to assess the contribution of each component: (a) w/o understanding, which removes the prompt-based GNN; (b) w/o inference, which eliminates multi-instruction fine-tuning; and (c) w/o both, which directly prompts LLMs using raw textual attributes. Figure 6 reports the node classification accuracy, where our full model consistently outperforms all variants. Specifically, the performance drop in w/o understanding highlights the importance of replicating GNN workflow to preserve both semantic and structural signals. The sharp decline in w/o inference further emphasizes the necessity of LLM fine-tuning for incorporating domain-specific knowledge. Finally, w/o both yields the worst results, underscoring the complementarity between the graph understanding and inference modules.
Figure 5: Ablation studies on node classification.
Figure 6: Impact of layer numbers.
Table 8: Quantitative graph vocabulary analysis on coverage and entropy.
Metric Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS ALL
Coverage (%) 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
Entropy 9.1116 9.4561 8.4643 10.7464 10.3752 11.3272 9.3825 12.2761
@@
Table 9: Efficiency analysis in terms of API cost (USD), number of tokens (M), training time (min), and power consumption (kWh).
@@
Efficiency Analysis (RQ5). We evaluate the computational and monetary cost of PromptGFM. As summarized in Table 9, while PromptGFM incurs an initial API cost during the graph understanding phase, this overhead is one-time and remains manageable with modern infrastructure. Specifically, each graph is processed only once to generate language-based node representations, which can be cached and reused throughout the inference pipeline. As a result, when additional graphs are introduced for training, there is no need to repeatedly call the LLM API to regenerate node representations. In contrast, traditional end-to-end models often require retraining from scratch when the graph structure changes, leading to higher cumulative cost. These characteristics make the computational overhead acceptable and amortizable, especially in dynamic or large-scale graph scenarios, where repeated retraining would otherwise be prohibitively expensive.
Prompt-based GNN Analysis (RQ6). We examine how the number of layers in our prompt-based GNN affects performance. As shown in Figure 6, increasing the number of layers initially improves results by capturing broader context and higher-order dependencies. However, beyond a certain depth, performance plateaus or declines due to over-smoothing, a trend consistent with traditional GNNs. Optimal performance is achieved with 3 layers on Cora and 2 layers on Citeseer. This suggests that textual representations can be propagated over the graph similarly to numerical embeddings, effectively capturing both semantics and structure. We further provide a case study on layer-wise representations in Appendix H.1, and compare language-based IDs with citation keywords in Appendix H.2. Additionally, we investigate permutation sensitivity of the prompts in Appendix G.
6 Conclusion
@@
APPENDIX
This appendix contains additional details for our paper, which is organized as follows:
$\bullet$ §A provides Data Descriptions used in our experiments.
$\bullet$ §B illustrates more details about Baselines employed for comparison.
$\bullet$ §C shows comparison between Embedding-based and Prompt-based GNNs.
$\bullet$ §D analyzes Prompt Design in our PromptGFM framework.
$\bullet$ §E shows more Implementation Details.
$\bullet$ §F provides a Prompt Length Limitation Discussion.
$\bullet$ §G investigates Permutation Sensitivity in prompt-based GNNs.
$\bullet$ §H presents the Case Study to offer deeper insights of our research.
Appendix A Data Descriptions
Dataset Domain #Nodes #Edges #Classes Raw Text
Cora Computer Science 2708 10858 7 paper titles and abstracts
Citeseer Computer Science 3327 9464 6 paper titles and abstracts
Ogbn-arxiv Computer Science 169343 2332486 40 paper titles and abstracts
WikiCS Computer Science 11701 431726 10 wikipedia entry names and contents
Photo E-commerce 48362 873782 12 item titles and reviews
History E-commerce 41551 503180 12 item titles and descriptions
Pubmed Biomedical 19717 88648 3 paper titles and abstracts
Table 10: Statistics of seven benchmarking datasets from three domains.
We utilize seven public benchmarking datasets from three domains to evaluate our framework, and the statistics of these datasets are illustrated in Table 10. For consistency, all graphs are treated as undirected. Below, we provide detailed descriptions of each dataset.
A.1 Computer Science Datasets
$\bullet$ Cora [32]. The dataset includes a citation network consisting of 2,708 scientific publications in the field of machine learning, categorized into 7 classes based on their research topics. Nodes represent individual papers, and edges denote citation links between them, totaling 10,858 connections. Each paper is associated with its title and abstract.
$\bullet$ Citeseer [13]. This dataset introduces a citation network comprising 3,327 scientific publications, categorized into 6 classes, including Agents, Artificial Intelligence, Database, Information Retrieval, Machine Learning, and Human-Computer Interaction. Nodes correspond to documents, and edges represent citation relationships between them, amounting to 9,464 links.
$\bullet$ obgn-arxiv [18]. The ogbn-arxiv dataset is part of the Open Graph Benchmark and consists of a directed citation graph of 169,343 arXiv papers, categorized into 40 subject areas. Nodes represent individual papers, and edges indicate citation relationships, totaling over 2.3 million connections. Each paper includes textual data from its title and abstract.
$\bullet$ WikiCS [33]. The WikiCS dataset is a benchmark dataset derived from Wikipedia, designed for evaluating GNNs. It comprises a citation network with 11,701 nodes representing computer science articles and 431,726 edges corresponding to hyperlinks between them. The dataset features 10 classes corresponding to branches of computer science, with very high connectivity.
A.2 E-commerce Datasets
$\bullet$ History [57]. The History dataset, extracted from the Amazon-Books dataset, includes items categorized under the second-level label "History". It comprises 41,511 nodes and 503,180 edges, where each node represents a book, and edges indicate frequent co-purchases or co-views between books. This dataset incorporates the title and description of each book as text attributes of the node. The classification task involves assigning books to 12 distinct categories.
$\bullet$ Photo [57]. The Photo dataset, derived from the Amazon-Electronics dataset, consists of 48,362 nodes and 873,782 edges, forming a dense network that reflects user purchasing behavior. Each node represents an electronic product, while edges indicate frequent co-purchases or co-views between items. For textual attributes, user reviews are incorporated, prioritizing the most upvoted review when available; otherwise, a randomly selected review is used. The classification task involves categorizing electronic products into 12 distinct categories.
A.3 Biomedical Datasets
$\bullet$ PubMed [42]. PubMed is a citation network of 19,717 scientific publications from the PubMed database pertaining to diabetes, classified into 3 classes: experimental induced diabetes, type 1 diabetes, and type 2 diabetes. Nodes are research papers, and edges signify citation links, amounting to 88,648 connections. This dataset is used for large-scale graph representation learning and evaluating algorithms in the biomedical domain.
Appendix B Baselines
@@
B.1 Graph-agnostic methods.
@@
$\bullet$ MLP. This method adopts a multi-layer perceptron to learn low-dimensional embeddings for each node. In our work, we randomly initialize node embeddings without textual attributes.
B.2 GNN-based methods.
$\bullet$ GCN  [23]. This model introduces a neural network architecture that generalizes convolution operations to graph-structured data, enabling effective semi-supervised learning by aggregating feature information from a node’s local neighborhood.
$\bullet$ GAT  [46]. This method incorporates attention mechanisms into graph neural networks, allowing nodes to assign different importance weights to their neighbors during feature aggregation, which enhances performance by focusing on the most relevant connections.
$\bullet$ SAGE  [16]. GraphSAGE is an inductive representation learning framework on large graphs; it generates node embeddings by sampling and aggregating features from a node’s local neighborhood, facilitating generalization to unseen nodes or graphs.
$\bullet$ ReVGNN  [24]. This method includes a recurrent graph neural network tailored for dynamic graphs, capturing temporal dependencies by updating node representations as events occur over time, which is crucial for modeling evolving graph structures.
@@
$\bullet$ SGFormer  [54]. This work introduces a transformer-based architecture designed for graph data, integrating spectral graph theory into the transformer framework. It aims to capture both local and global graph structures efficiently by incorporating spectral filters, enhancing the model’s ability to learn complex graph representations.
@@
$\bullet$ NodeFormer  [53]. This framework presents a scalable graph transformer model that utilizes a randomized attention mechanism to approximate full attention on graphs. By reducing computational complexity, it enables efficient learning on large-scale graphs while preserving the expressiveness of transformer architectures.
B.3 LM-only methods.
$\bullet$ BERT  [10]. This work introduces a deep learning model that understands language context by processing text bidirectionally. It is pre-trained using masked language modeling (MLM) and next sentence prediction (NSP) to learn rich linguistic features. BERT can be fine-tuned for various NLP tasks, achieving state-of-the-art performance.
$\bullet$ Sentence-BERT  [40]. This model is a modification of BERT designed for sentence embeddings, allowing efficient comparison of semantic similarity. It uses a siamese or triplet network structure to generate fixed-size vector representations, making tasks like sentence similarity, clustering, and retrieval significantly faster. Unlike BERT, which requires computationally expensive cross-encoding, Sentence-BERT enables quick and effective comparisons using cosine similarity.
$\bullet$ LLaMA3  [27]. LLaMA3-8B is an open-weight transformer model with 8 billion parameters, part of Meta’s third-generation LLaMA series. It is trained on 15T tokens from a diverse mix of web, code, and academic data, using a decoder-only architecture with a context window of 8,192 tokens. LLaMA3-8B achieves strong performance on a wide range of benchmarks, particularly in reasoning, multilingual tasks, and instruction following. Compared to earlier versions, it benefits from improved training stability, data quality, and scaling efficiency, making it a competitive foundation model for downstream adaptation and alignment.
B.4 GNN-LLM integration methods.
$\bullet$ LLaGA  [6]. This model effectively integrates LLM capabilities to handle the complexities of graph-structured data. It transforms graph nodes into structure-aware sequences and maps them into token embedding space using a specialized projector. LLaGA excels in generalization and interpretability, performing strongly across various datasets and tasks. It also supports zero-shot learning, making it highly adaptable for unseen datasets.
$\bullet$ OFA  [27]. This paper proposes a framework that handles various graph classification tasks across different domains using a single model. It introduces the nodes-of-interest (NOI) subgraph mechanism to standardize different tasks with a single task representation. Additionally, a novel graph prompting paradigm to leverage in-context learning and apply the same architecture across diverse graph classification tasks, achieving generalization across multiple domains.
@@
$\bullet$ GraphPrompter  [29]. This work introduces a novel framework designed to align graph with LLMs via soft prompts. Specifically, it adopts GNNs to capture graph structure and leverages an LLM to interpret the textual information at the node level. By prompt tuning, this approach demonstrates the potential of LLMs to effectively interpret graph structures, combining both semantic and structural insights for improved graph learning tasks.
@@
$\bullet$ ENGINE  [62]. This paper proposes a parameter- and memory-efficient fine-tuning method for textual graphs by using LLMs as encoders. It combines the LLMs and GNNs through a tunable GNN-based side structure, called G-Ladder, alongside each LLM layer, effectively reducing training costs without compromising performance.
Figure 7: Graph understanding module via prompt-based GNNs. We prompt LLMs to achieve fine-grained reproduction of traditional GNN workflow, refining verbose textual representations into concise yet meaningful ones. In the prompt, neighbor sampling (see Equation 3) is highlighted in darkred, the aggregation-update mechanism (see Equation 4) in blue, and the optimization in red.
Appendix C Embedding-based GNNs vs. Prompt-based GNNs
Figure 7 illustrates the details of the graph understanding module, with the prompt-based GNN as its core component. To ensure clarity, we systematically compare traditional embedding-based GNNs with our proposed prompt-based GNN, emphasizing its advantages.
$\bullet$ Input and output. In the embedding-based GNN framework, for each node, structure-less embeddings are progressively refined to structure-rich embeddings, whereas verbose textual sequences are gradually converted to concise textual sequences in our prompt-based GNN.
$\bullet$ Message passing. The multi-layer embedding updates are mirrored by multi-round LLM calls in text space, both of which progressively refining the representations in different spaces.
$\bullet$ Neighbor sampling. The neighbor sampling operation used to reduce computational load in traditional embedding-based GNNs is analogous to the selected one-hop neighbor descriptions employed to address prompt length limitations in prompt-based GNNs.
$\bullet$ Aggregation-update mechanism. Embedding-based GNNs use predefined operators (e.g., mean aggregator, weighted aggregator, or LSTM aggregator) to achieve message passing in the embedding space, while prompt-based GNNs use straightforward prompts to guide LLMs in executing the process more flexibly without predefined rules.
$\bullet$ Optimization. In prompt-based GNNs, we use heuristic prompts at each layer to reflect the key idea of contrastive loss. These cumulative layer-by-layer prompts are comparable to the layer-wise loss combination, formally as mean pooling, commonly seen in embedding-based GNNs.
In conclusion, our approach enables LLMs to replicate the finest GNN workflow within the text space, fostering seamless graph-text alignment. By bridging structural and semantic information through language-based representations, this framework introduces a novel paradigm for GNN-LLM integration. It not only enhances the interpretability and transferability of graph learning but also unlocks new possibilities for leveraging LLMs in graph-related tasks without relying on numerical embeddings.
Appendix D Prompt Design
In this section, we provide the templates of prompts in our PromptGFM framework, including prompt-based GNNs and task-oriented prompts.
Prompt for node initialization. The title of the paper is <the title of the paper>, the abstract of the paper is <the abstract of the paper>. Please summarize the paper.
/*** This prompt varies depending on the dataset. The instance above is designed for citation datasets. ***/
Prompt for each layer of GNN replication. Given the central node <l-th round textual representation of the central node>. The selected one-hop neighbors are [< l-th round of node #1>, <l-th round node #2>, ... ,<l-th round node #N>]. Please aggregate neighbor nodes and update a concise yet meaningful representation for the central node. Note connected nodes should share similar semantics and vice versa.
Prompt for node classification. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Which category should <the language-based ID of the central node> be classified as ?
Prompt for discriminative link prediction. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Among <the language-based ID of the central node> and <the language-based ID of its negative sampling node>, which node will be connected to <the language-based ID of the central node> ?
Prompt for generative link prediction. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Which node should be connected to <the language-based ID of the central node>?
Appendix E Implementation Details

### v3

@@ 5.3. Cross-graph and Cross-task Transfer (RQ2)
ST-NC 91.72 84.49 92.83
LP zero-shot 67.28 69.22 73.59
@@
5.4. Diverse Pretraining Tasks (RQ3)
A robust foundation model should support a rich set of pre-training tasks, as task diversity is expected to enhance generalization on downstream applications. To this end, we extend beyond the NC and LP tasks by introducing two additional pre-training objectives: (a) Node Discrimination (ND): Given three nodes [a], [b], [c], exactly one belongs to a different category. Which one is it?. and (b) Link Discrimination (LD): Given a central node [X], which candidate node [a], [b], [c] is not connected to [X]?, where [x] denotes the language-based ID of a node. Besides NC and LP, we incorporate ND and LD individually or jointly into multi-task fine-tuning, and evaluate their effects on NC performance under otherwise identical settings. As shown in Table 8, we observe that adding either ND or LD improves both accuracy and M-F1, while combining the two yields the best performance. These results validate that expanding the pre-training task set leads to richer representations. More importantly, they demonstrate that PromptGFM does not merely memorize task patterns, but leverages its unified language-based interface to generalize across tasks and exhibit emergent capabilities, a defining property of foundation models.
Table 8. The impact of incorporating Node Discrimination (+ND) and Link Discrimination (+LD) as pre-training tasks on NC performance.
Dataset +ND +LD Accuracy (%) Macro-F1 (%)
Citeseer 84.49 82.31
85.67 ($\uparrow$1.18) 83.22 ($\uparrow$0.91)
84.73 ($\uparrow$0.24) 82.65 ($\uparrow$0.34)
86.02 ($\uparrow$1.53) 83.77 ($\uparrow$1.46)
PubMed 90.67 91.82
92.53 ($\uparrow$1.86) 92.84 ($\uparrow$1.02)
91.12 ($\uparrow$0.45) 92.03 ($\uparrow$0.21)
93.06 ($\uparrow$2.39) 93.27 ($\uparrow$1.45)
Table 9. Permutation sensitivity of prompt-based GNNs.
Variant Cora Citeseer
shuffle nodes 90.64 $\pm$1.12 83.79 $\pm$ 0.89
shuffle tokens 90.55 $\pm$1.03 84.28 $\pm$1.06
Table 10. Layer-wise mean cosine similarity of node embeddings on Cora, with margin reflecting the gap between connected and disconnected pairs.
Connected Pairs Disconnected Pairs Margin
Layer 1 0.171 0.036 0.135
Layer 2 0.223 0.032 0.191
Layer 3 0.308 0.034 0.274
5.5. Prompt-based GNN Analysis (RQ4)
5.5.1. Permutation Sensitivity
We evaluate the sensitivity of our prompt-based GNNs (using Flan-T5) to input order by evaluating two variants: (a) shuffle nodes, which randomizes the order of neighbors, and (b) shuffle tokens, which injects position-specific tokens to nodes and shuffle them instead. As illustrated in Table 9, regardless of the shuffle method, both variants result in only minor fluctuations in node classification accuracy, indicating that prompt-based GNNs with the subsequent fine-tuning are robust to input permutation and maintain manageable and stable performance.
Table 11. Quantitative graph vocabulary analysis.
Metric Cora Citeseer PubMed Ogbn-arxiv
Coverage (%) 100.00 100.00 100.00 100.00
Entropy 9.1116 9.4561 8.4643 10.7464
Metric History Photo WikiCS ALL
Coverage (%) 100.00 100.00 100.00 100.00
Entropy 10.3752 11.3272 9.3825 12.2761
Figure 5. Ablation studies on node classification.
Figure 6. Impact of layer numbers.
5.5.2. Representation Discrimination over Layers
To validate that our language-based graph vocabulary effectively preserves structural proximity, we conduct an analysis by encoding the final language-based IDs with a pretrained Sentence-BERT encoder and computing the cosine similarity between node pairs. The results on the Cora dataset are presented in Table 10. We observe that connected nodes exhibit a significantly higher mean similarity than disconnected nodes ($p<10^{-5}$). Furthermore, the margin between the two groups grows steadily as the number of GNN layers increases, indicating stronger structural discrimination with deeper propagation. This trend highlights that our prompt-based GNN captures meaningful semantics with stacking aggregation-update layers, thereby demonstrating its theoretical equivalence in representational power to traditional embedding-based GNNs.
5.5.3. Performance over GNN Layers.
We investigate the impact of layer depth in our prompt-based GNN. As shown in Figure 6, increasing the number of layers initially improves results by capturing broader context and higher-order dependencies. However, beyond a certain depth, performance plateaus or declines due to over-smoothing, a phenomenon also observed in traditional GNNs. The optimal depth is achieved with three layers on Cora and two layers on Citeseer. These results suggest that textual representations can be propagated across the graph in much the same way as numerical embeddings, effectively encoding both semantic and structural information. We further provide a case study on layer-wise representations in Appendix C.3.
5.6. Exploration Studies
5.6.1. Ablation Studies.
@@ 5.6. Exploration Studies
We construct three ablated variants to assess the contribution of each component: (a) w/o understanding, which removes the prompt-based GNN; (b) w/o inference, which eliminates multi-instruction fine-tuning; and (c) w/o both, which directly prompts LLMs using raw textual attributes. Figure 6 reports the node classification accuracy, where our full model consistently outperforms all variants. Specifically, the performance drop in w/o understanding highlights the importance of replicating GNN workflow to preserve both semantic and structural signals. The sharp decline in w/o inference further emphasizes the necessity of LLM fine-tuning for incorporating domain-specific knowledge. Finally, w/o both yields the worst results, underscoring the complementarity between the graph understanding and inference modules.
5.6.2. Vocabulary Analysis.
We evaluate the quality of our graph vocabulary using two metrics in Table 11. (a) Coverage. We compute the proportion of nodes successfully matched to IDs. Since the graph vocabulary is grounded in LLM’s native language, any node’s semantics can, in principle, be expressed through a combination of natural language tokens. Thus, every node is assigned a valid ID, confirming full coverage. (b) Uniqueness. We measure the entropy of token distributions across nodes within both individual and merged graphs. Higher entropy indicates more diverse and fine-grained token usage, indicating stronger uniqueness and representational capacity. Notably, the merged vocabulary (“ALL”) achieves the highest entropy (12.2761), showing that integrating multiple graphs improves vocabulary expressiveness and discriminability, which reinforces the necessity of building GFMs.
5.6.3. Efficiency Analysis.
We evaluate the computational and monetary cost of PromptGFM, as summarized in Table 12.In the graph understanding module, the API or token cost grows with the graph size, but this constitutes a one-time, graph-specific overhead that can be cached and reused throughout the inference pipeline (denoted as ”n/a”). During inference, as more graphs are introduced and the overall vocabulary expands, we observe that time and energy consumption increase in an approximately linear manner. These results align with our discussion in Section 4.3 and and demonstrate PromptGFM’s efficiency as a scalable GFM.
@@
Table 12. Efficiency analysis in terms of API cost (USD), number of tokens (M), inference time (min), and power consumption (kWh).
@@
6. Conclusion
@@
Appendix A Experimental Setup
A.1. Data Descriptions
We use seven public benchmarking datasets from three domains to evaluate PromptGFM, and the statistics are illustrated in Table 13.
A.2. Baselines
@@
A.2.1. Graph-agnostic methods.
@@
$\bullet$ MLP. This method adopts a multi-layer perceptron to learn low-dimensional embeddings for each node.
A.2.2. GNN-based methods.
$\bullet$ GCN  (Kipf and Welling, 2017). This model generalizes convolutional operations to graphs for semi-supervised learning.
$\bullet$ GAT  (Velickovic et al., 2017). This method integrates attention mechanisms to weigh neighbor features during aggregation.
$\bullet$ SAGE  (Hamilton et al., 2017). GraphSAGE is an inductive framework which enables generalization to unseen data.
$\bullet$ ReVGNN  (Li et al., 2021). This method includes a recurrent GNN that updates node representations as the graph evolves over time.
@@
$\bullet$ SGFormer  (Wu et al., 2023). This work introduces a Transformer-based architecture that integrates spectral filters to capture both local and global graph structures.
@@
$\bullet$ NodeFormer  (Wu et al., 2022). This framework presents a scalable graph Transformer that utilizes a randomized attention mechanism, enabling efficient training on large-scale graphs.
A.2.3. LM-only methods.
$\bullet$ BERT  (Devlin et al., 2019). This language model learns bidirectional context through pre-training with masked language modeling (MLM) and next sentence prediction (NSP).
$\bullet$ Sentence-BERT  (Reimers and Gurevych, 2019). This model is a modification of BERT that uses a siamese network to generate fixed-size sentence embeddings.
$\bullet$ LLaMA3  (Liu et al., 2024a). This is an open-weight, decoder-only Transformer with 8 billion parameters, which achieves strong performance on a wide range of benchmarks, particularly in reasoning, multilingual tasks, and instruction following.
A.2.4. GNN-LLM integration methods.
$\bullet$ LLaGA  (Chen et al., 2024c). This model integrates LLMs for graph data by transforming nodes into structure-aware sequences and mapping them into the token embedding space with a projector.
$\bullet$ OFA  (Liu et al., 2024a). This framework handles various graph classification tasks across different domains using a single model.
@@
$\bullet$ GraphPrompter  (Liu et al., 2024b). This framework aligns graphs with LLMs via soft prompts, including a GNN to capture structure and an LLM to interpret node text, combining semantic and structural insights through prompt tuning.
@@
$\bullet$ ENGINE  (Zhu et al., 2024). This paper presents a parameter- and memory-efficient fine-tuning method, which combines a tunable GNN-based side structure alongside each LLM layer.
A.3. Implementation Details